OpenStudy (anonymous):

A wooden artifact from an ancient tomb contains 30 percent of the carbon-14 that is present in living trees. How long ago was the artifact made? (The half-life of carbon-14 is 5730 years.) Your answer is ? years.

OpenStudy (anonymous):

\[\frac{1}{2}^{\frac{t}{5730}} = \frac{3}{10}\] Solve for t.